# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ 3.1. Random fields on surfaces
where $\alpha>d/2$. By selecting $D_{ij}(x)=\delta_{ij}$ and $V(x)=\kappa^{2}$ with $\kappa>0$ one recovers the classical, stationary Whittle–Matérn fields studied in various settings in for instance Cox and Kirchner (2020); Bolin et al. (2020); Bonito et al. (2024); Lindgren et al. (2011); Jansson et al. (2022); Whittle (1963). In Figures 4(a) and 4(b) we show this case with $\kappa^{2}=10$ for a rougher field with $\alpha=0.55$ and a smoother one with $\alpha=1.5$, respectively.
@@ 3.2. Approximation of random fields with surface finite elements
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. Following the approach presented by Lang and Pereira (2023), the field $\mathcal{Z}$ is approximated by an expansion similar to that of Equation 27, but involving only quantities defined on the polyhedral surface $\mathcal{M}_{h}$. More precisely, we define the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ of the field $\mathcal{Z}$ by the relation
@@
where $\{W_{i}^{h}\}_{1\leq i\leq N_{h}}$ is a sequence of independent standard Gaussian random variables, whose precise definition is clarified later in this section, and $\{(\Lambda_{i}^{h},E_{i}^{h})\}_{1\leq i\leq N_{h}}$ are the eigenpairs of $\mathsf{L}_{h}$ introduced in Section 2.2. Then, as proven in Lang and Pereira (2023, Theorem 3.4), $\mathsf{Z}_{h}$ can be decomposed in the nodal basis $\{\psi_{i}\}_{1\leq i\leq N_{h}}$ of $S_{h}$ as
@@
Note that sampling the weights $(Z_{1},\dots,Z_{N_{h}})$, and therefore the field $\mathsf{Z}_{h}$, using directly the expression of their covariance matrix requires in practice to fully diagonalize $\bm{S}$ (since Equation 29 involves a function of a matrix). Such an operation would result in a prohibitive computational cost (of order $\mathcal{O}(N_{h}^{3})$ operations). To avoid this cost, we use the Chebyshev trick proposed by Lang and Pereira (2023, Section 4), and approximate $\mathsf{Z}_{h}$ by the field $\mathsf{Z}_{h,M}$ defined by
@@
where $P_{\gamma,M}$ is a Chebyshev polynomial approximation of degree $M\in\amsmathbb{N}$ of $\gamma$ over an interval $[\lambda_{\min},\lambda_{\max}]$ containing all the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ of $\mathsf{L}_{h}$ (which we recall, coincide with the eigenvalues of $\bm{S}$). Such an interval can be obtained as follows. On the one hand, one can take $\lambda_{\min}=V_{-}$. On the other hand, following Lang and Pereira (2023), a candidate for $\lambda_{\max}$ is obtained by applying the Gershgorin circle theorem to $\bm{S}$.
Since $\mathsf{Z}_{h,M}$ is defined by just replacing the power spectral density $\gamma$ by the polynomial  $P_{\gamma,M}$, its expansion $\mathsf{Z}_{h,M}$ into the nodal basis,
$\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}Z_{i}^{(M)}\psi_{i},$
@@
$\begin{pmatrix}Z_{1}^{(M)}\\ \vdots\\ Z_{N_{h}}^{(M)}\end{pmatrix}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}P_{\gamma,M}(\bm{S})\begin{pmatrix}\varepsilon_{1}\\ \vdots\\ \varepsilon_{N_{h}}\end{pmatrix}$
where $\varepsilon_{1},\dots,\varepsilon_{N_{h}}$ are independent standard Gaussian random variables, and the matrix-vector product by $P_{\gamma,M}(\bm{S})$ can be computed iteratively while just requiring products between $\bm{S}$ and vectors. In the next two subsections, we provide error estimates quantifying the error between our target random field $\mathcal{Z}$ and its successive SFEM and polynomial approximations by $\mathsf{Z}_{h}$ and $\mathsf{Z}_{h,M}$.
@@ 3.3. Error analysis of the SFEM discretization
Let $\gamma$ be an $\alpha$-power spectral density with $\alpha>d/4$. Then, there exists $h_{0}>0$, such that for any $h<h_{0}$ the strong approximation error of the random field $\mathcal{Z}$ by its discretization $\mathsf{Z}_{h}^{\ell}$ satisfies the bound
(33) $\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
where $K_{\alpha}(h)=1$ if $\alpha-d/4<2$, $K_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$, and $K_{\alpha}(h)=|\log h|^{(d-1)/2}$ otherwise.
@@
By injecting the representation (26) of $\mathcal{W}$ in the definition of $\xi_{j}$, we get that $\widetilde{\mathcal{W}}$ can itself be formally represented by $\widetilde{\mathcal{W}}=\sum_{k=1}^{\infty}W_{k}P_{h}e_{k}=P_{h}\mathcal{W}$. This explains why we refer to it as a projected white noise.
@@
(36) $\mathsf{Z}_{h}=\gamma(\mathsf{L}_{h})\mathsf{W}$
@@
$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h}$
@@
where $\beta_{j}=(\mathsf{W},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma^{1/2}\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},\sigma^{-1/2}E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$, and the same computations as before yield that $\beta_{j}=\big{(}\mathcal{W},(\sigma^{-1/2}E_{j}^{h})^{\ell}\big{)}_{L^{2}(\mathcal{M})}$. Hence, we can conclude this time that for any $1\leq i,j\leq N_{h}$, $\beta_{j}$ is also normally distributed with mean 00, and that
@@
$\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}$ $\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\tilde{\mathcal{Z}}_{h}-\widehat{\mathsf{Z}}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\widehat{\mathsf{Z}}_{h}^{\ell}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
$\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}{-}\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}+\big{\|}\widehat{\mathsf{Z}}_{h}-\mathsf{Z}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
@@
It holds that
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{\min\{\alpha-d/4;2\}},$
where $C_{\alpha}(h)=|\log h|$ if $\alpha-d/4=2$ and $C_{\alpha}(h)=1$, otherwise.
@@
Our aim is to bound the error between $\tilde{\mathcal{Z}}_{h}$ and $\mathcal{Z}$, where we remark in particular that $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W}$ and $\mathcal{Z}_{h}=\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}$ (cf. Remark 3.2). Let $\eta=\alpha+d/4>d/2$, and let $\zeta=\alpha-d/4>0$, so that $\alpha=(\eta+\zeta)/2$. Further, define the function $\tilde{\gamma}(x)=\gamma(x)x^{\eta/2}$. Note that, since $\gamma$ is an $\alpha$-power spectral density and since $\tilde{\gamma}$ decays as
$|\tilde{\gamma}(x)|\lesssim|x|^{-\left(\alpha-\eta/2\right)}=|x|^{-\zeta/2},$
$\tilde{\gamma}$ is a $(\zeta/2)$-power spectral density. Now, by definition of $\tilde{\gamma}$,
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $=\|\tilde{\gamma}(\mathcal{L})\mathcal{L}^{-\eta/2}\mathcal{W}-\tilde{\gamma}(\mathcal{L}_{h})\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
so that the triangle inequality yields
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq E_{1}+E_{2},$

### v2

@@ 3.1. Random fields on surfaces
where $\alpha>d/2$. By selecting $D_{ij}(x)=\delta_{ij}$ and $V(x)=\kappa^{2}$ with $\kappa>0$ one recovers the classical, stationary Whittle–Matérn fields studied in various settings in for instance [9, 3, 6, 21, 18, 27]. In Figures 4(a) and 4(b) we show this case with $\kappa^{2}=10$ for a rougher field with $\alpha=0.55$ and a smoother one with $\alpha=1.5$, respectively.
@@
Finally, we remark that other types of amplitude spectral densities can be considered. For instance, to generate Figure 1(c), we used an amplitude spectral density of the form $\gamma(\lambda)=\sin(\lambda)\lambda^{-0.6}$.
@@ 3.2. Approximation of random fields with surface finite elements
Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$. Following the approach presented by [19], the field $\mathcal{Z}$ is approximated by an expansion similar to that of Equation 29, but involving only quantities defined on the polyhedral surface $\mathcal{M}_{h}$. More precisely, we define the SFEM–Galerkin approximation $\mathsf{Z}_{h}$ of the field $\mathcal{Z}$ by the relation
@@
where $\{W_{i}^{h}\}_{1\leq i\leq N_{h}}$ is a sequence of independent standard Gaussian random variables, whose precise definition is clarified later in this section, and $\{(\Lambda_{i}^{h},E_{i}^{h})\}_{1\leq i\leq N_{h}}$ are the eigenpairs of $\mathsf{L}_{h}$ introduced in Section 2.2. Then, as proven in [19, Theorem 3.4], $\mathsf{Z}_{h}$ can be decomposed in the nodal basis $\{\psi_{i}\}_{1\leq i\leq N_{h}}$ of $S_{h}$ as
@@
Note that sampling the weights $(Z_{1},\dots,Z_{N_{h}})$, and therefore the field $\mathsf{Z}_{h}$, using directly the expression of their covariance matrix requires in practice to fully diagonalize $\bm{S}$ (since Equation 31 involves a function of a matrix). Such an operation would result in a prohibitive computational cost (of order $\mathcal{O}(N_{h}^{3})$ operations). To avoid this cost, we use the Chebyshev trick proposed by [19, Section 4], and approximate $\mathsf{Z}_{h}$ by the field $\mathsf{Z}_{h,M}$ defined by
@@
where $P_{\gamma,M}$ is a Chebyshev polynomial approximation of degree $M\in\amsmathbb{N}$ of $\gamma$ over an interval $[\lambda_{\min},\lambda_{\max}]$ containing all the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ of $\mathsf{L}_{h}$ (which we recall, coincide with the eigenvalues of $\bm{S}$). Such an interval can be obtained as follows. On the one hand, one can take $\lambda_{\min}=V_{-}$. On the other hand, following [19], a candidate for $\lambda_{\max}$ is obtained by applying the Gershgorin circle theorem to $\bm{S}$.
The field $\mathsf{Z}_{h,M}$, called Galerkin–Chebyshev approximation of $\mathcal{Z}$, is in essence defined by just replacing the amplitude spectral density $\gamma$ by the polynomial  $P_{\gamma,M}$ in the definition of SFEM–Galerkin approximation $\mathsf{Z}_{h}$. The expansion of $\mathsf{Z}_{h,M}$ into the nodal basis,
(34) $\mathsf{Z}_{h,M}=\sum_{i=1}^{N_{h}}Z_{i}^{(M)}\psi_{i},$
@@
(36) $\begin{pmatrix}Z_{1}^{(M)}\\ \vdots\\ Z_{N_{h}}^{(M)}\end{pmatrix}=\big{(}\sqrt{\bm{C}}\big{)}^{-T}P_{\gamma,M}(\bm{S})\bm{W}$
where $\bm{W}=(w_{1},\dots,w_{N_{h}})^{T}$ is a vector of independent standard Gaussian random variables, and the matrix-vector product by $P_{\gamma,M}(\bm{S})$ can be computed iteratively while just requiring products between $\bm{S}$ and vectors. In the next two subsections, we provide error estimates quantifying the error between our target random field $\mathcal{Z}$ and its successive SFEM and polynomial approximations by $\mathsf{Z}_{h}$ and $\mathsf{Z}_{h,M}$.
@@ 3.3. Error analysis of the SFEM discretization
Let $\gamma$ be an $\alpha$-amplitude spectral density with $\alpha>d/4$. Then, for any $h\in(0,h_{0})$, the strong approximation error of the random field $\mathcal{Z}$ by its discretization $\mathsf{Z}_{h}^{\ell}$ satisfies the bound
(37) $\|\mathcal{Z}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
@@
By injecting the representation (28) of $\mathcal{W}$ in the definition of $\xi_{j}$, we get that $\widetilde{\mathcal{W}}$ can itself be formally represented by $\widetilde{\mathcal{W}}=\sum_{k=1}^{\infty}W_{k}P_{h}e_{k}=P_{h}\mathcal{W}$. This explains why we refer to it as a projected white noise. In particular, note that $\tilde{\mathcal{W}}\in L^{2}(\Omega;L^{2}(\mathcal{M}))$ and $\|\tilde{\mathcal{W}}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{j=1}^{N_{h}}\xi_{j}e_{j}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=N_{h}<\infty$.
@@
(40) $\mathsf{Z}_{h}=\gamma(\mathsf{L}_{h})\mathsf{W},$
@@
$\widehat{\mathsf{W}}=\sum_{j=1}^{N_{h}}\alpha_{j}E_{j}^{h},$
@@
where $\beta_{j}=(\mathsf{W},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma^{1/2}\widetilde{\mathcal{W}}^{-\ell},E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma\widetilde{\mathcal{W}}^{-\ell},\sigma^{-1/2}E_{j}^{h})_{L^{2}(\mathcal{M}_{h})}$, and the same computations as before yield that $\beta_{j}=\big{(}\mathcal{W},(\sigma^{-1/2}E_{j}^{h})^{\ell}\big{)}_{L^{2}(\mathcal{M})}$. Hence, we conclude this time that for any $1\leq i,j\leq N_{h}$, $\beta_{j}$ is also normally distributed with mean 00, and that
@@
[…]
$\quad\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\tilde{\mathcal{Z}}_{h}-\widehat{\mathsf{Z}}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}+\|\widehat{\mathsf{Z}}_{h}^{\ell}-\mathsf{Z}_{h}^{\ell}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$
$\quad\lesssim\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}{+}\big{\|}\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}{-}\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}{+}\big{\|}\widehat{\mathsf{Z}}_{h}-\mathsf{Z}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
@@
Let $h\in(0,h_{0})$. It holds that
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}},$
where $C_{\alpha}(h)=|\log h|$ if $d/4<\alpha\leq 1$, $C_{\alpha}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2, and $C_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$.
@@
Our aim is to bound the error between $\tilde{\mathcal{Z}}_{h}$ and $\mathcal{Z}$, where we remark in particular that $\mathcal{Z}=\gamma(\mathcal{L})\mathcal{W}$ and $\mathcal{Z}_{h}=\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}$ (cf. Remark 3.2). Recall that we consider $h\in(0,h_{0})$ (meaning in particular that $|\log h|>1$).
Similarly to the proof of [6, Theorem 5.2], we introduce $\varepsilon_{h}=\min\{d/2;\alpha-d/2\}/(2|\log h|)$, $\eta=2(d/4+\varepsilon_{h})$, and let $\zeta=2(\alpha-d/4-\varepsilon_{h})>0$, so that $\alpha=(\eta+\zeta)/2$. Note in particular that $\varepsilon_{h}>0$ since $d\in\{1,2\}$ and $\alpha>1>d/2$, and $\varepsilon_{h}<(\alpha-d/2)/2<\alpha-d/2<\alpha-d/4$. Besides, $\eta>d/2$ and $\zeta>0$. We distinguish two cases: if $\alpha\leq 1$, and if $\alpha>1$.
Case $\alpha\leq 1$
The proof follows the same approach as [6, Theorem 5.2]. Indeed, note that the triangle inequality yields
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}=\|\gamma(\mathcal{L})\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq S_{1}+S_{2},$